OpenStudy (anonymous):

true or false? 8/15 = 1/2

jimthompson5910 (jim_thompson5910):

Multiply top and bottom of 8/15 by 2 to get 16/30 on the left side Multiply top and bottom of 1/2 by 15 to get 15/30 on the right side So does 16/30 = 15/30 ? The answer is "no", so \[\Large \frac{8}{15}\neq \frac{1}{2}\]

jhonyy9 (jhonyy9):

false